### v1

# Discovering dense hydrogen solid at 1200K with deep variational free energy approach

Abstract
We perform deep variational free energy calculations to investigate the dense hydrogen system at 1200 K and high pressures. In this computational framework, neural networks are used to model the free energy through the proton Boltzmann distribution and the electron wavefunction. By directly minimizing the free energy, our results reveal the emergence of a crystalline order associated with the center of mass of hydrogen molecules at approximately 180 GPa. This transition from atomic liquid to a molecular solid is marked by discontinuities in both the pressure and thermal entropy. Additionally, we discuss the broader implications and limitations of these findings in the context of recent studies of dense hydrogen under similar conditions.
I Introduction
Dense hydrogen exhibits a diverse range of fascinating physical phenomena [1] despite being composed of the simplest element. At low temperatures, the system adopts a variety of solid-state phases, which have been observed in both experiments [2, 3, 4, 5] and simulations [6, 7]. At high temperatures, the system exists in a liquid phase, consisting of either molecular or atomic hydrogen [8, 9]. Understanding the transition between these two liquid states is essential for both planetary modeling and high-pressure physics [10, 11, 12, 13, 14].
In spite of extensive computational studies of the liquid-liquid transition in the hydrogen system, its underlying physics remains unclear. This is partly due to that traditional methods, such as density functional theory (DFT) or ab-initio molecular dynamics, are limited by their accuracy [15, 16, 17]. More accurate yet computationally expensive methods, such as quantum Monte Carlo, also face challenges in unbiasedly sampling the proton degrees of freedom [18, 19, 20, 21, 22]. Moreover, most of these approaches suffer from high computational costs, which limit simulations to relatively small systems or short timescales. Recent efforts based on machine-learned force fields [23] have enabled simulations at much larger sizes and longer timescales. However, the possible trade-off in accuracy during the supervised training [24] may limit their capacity to fully capture the complex behavior of the underlying physical processes. For example, the latest molecular dynamics studies based on the machine learning force field [25, 26] show qualitatively different results compared to earlier works [23]. In this context, it is noteworthy to mention recent findings on the solid-state phase around 1200 K [27, 28] obtained using machine learning force fields trained with quantum Monte Carlo (QMC) data, which sharply contrast with previous DFT calculations and experimental observations.
In order to investigate possible phases of dense hydrogen in this parameter regime, we perform a variational free energy calculation [29] to study the system at 1200 K. This computational method combines flow-based generative models and neural network-based wavefunctions to model the variational density matrix of the dense hydrogen system. By modeling the proton Boltzmann distribution and electron wavefunction using two deep neural networks and minimizing the free energy, we obtain an approximate solution for the system’s equation of state at the QMC level of accuracy, which is approximately 1 milli-Hartree per atom. Additionally, we incorporate physical knowledge into the model through a pretraining strategy based on proton ensembles and utilize a machine learning force field as the base distribution of the flow model. This computational approach is closely related to coupled electron-ion Monte Carlo [30] and Langevin dynamics [31] methods. The key difference is that we replace the Monte Carlo sampling of protons with a variational calculation, which provides direct access to the free energy and entropy of the system. All of these methods take into account many-electron correlations [32] and thermal effects, but exhibit different characteristics when applied to finite-temperature structure prediction [33]. Molecular dynamics or Monte Carlo simulations are unbiased methods for sampling the potential energy surface. However, they may face challenges in capturing transitions between different phases, which correspond to different modes in the probability distribution, thus hindering the discovery of novel structures. In contrast, the variational free energy calculation has the “mode seeking” behavior [34], which can lead to convergence at local minima in the free energy landscape. Fortunately, having access to the variational free energy provides a way to check the optimality of the obtained solution.
This paper is organized as follows: Section II introduces the computational framework of the deep variational free energy approach, which contains a flow model enhanced by machine learning force fields and neural network wavefunction. Section III presents benchmark results and analyzes the discovered solid-state phases in the dense hydrogen system. Finally, Section IV discusses the implications and limitations of the present study.
II Methods
II.1 Variational free energy framework
We consider a system of $N$ hydrogen atoms confined in a periodic simulation cell, with cell vectors $\bm{a}=(\bm{a}_{1},\bm{a}_{2},\bm{a}_{3})$ and reciprocal vectors $\bm{a}^{\ast}=(\bm{a}_{1}^{\ast},\bm{a}_{2}^{\ast},\bm{a}_{3}^{\ast})$. In our convention, $\bm{a}_{\mu}\cdot\bm{a}_{\nu}^{\ast}=\delta_{\mu\nu}$, where $\delta$ is the Kronecker delta. Throughout this work, the coordinates of the particles and the cell vectors are scaled by a dimensionless parameter $1/r_{s}$, which represents the length scale of the system. For a cubic simulation cell with side length $L$, $r_{s}=\left(3/4\pi N\right)^{1/3}L/a_{B}$, where $a_{B}$ is the Bohr radius. The proton and electron coordinates are denoted by $\bm{X}=\{\bm{x}_{I}\}$ and $\bm{R}=\{\bm{r}_{i}\}$, respectively. For convenience, we sometimes treat the protons and electrons on equal footing and introduce $\bm{\Xi}=\{\bm{\xi}_{\alpha}\}=\bm{R}\bigcup\bm{X}$ as a unified notation for the coordinates of both electrons and protons. Using this convention, the Hamiltonian of the system is written as
$H=-\frac{1}{2r_{s}^{2}}\sum_{i}\nabla_{\bm{r}_{i}}^{2}+\frac{1}{2r_{s}}\sum_{\alpha\beta\bm{n}}^{\ast}\frac{q_{\alpha}q_{\beta}}{|\bm{\xi}_{\alpha}-\bm{\xi}_{\beta}+\bm{n}|}\,,$ (1)
where $q_{\alpha}$ is the charge of particle $\alpha$, and $\bm{n}=n_{1}\bm{a}_{1}+n_{2}\bm{a}_{2}+n_{3}\bm{a}_{3}$, with $n_{\mu}\in\mathbb{Z}$ and $\mu=1,2,3$. The $\ast$” in the summation indicates that if $\bm{n}=0$, the terms with $\alpha=\beta$ are omitted. We will focus on spin unpolarized systems with $N/2$ spin-up and $N/2$ spin-down electrons in this work.
Figure 1: The flowchart of pretraining and finetuning of the hydrogen variational density matrix. A force field provides both the proton configurations for pretraining of the wavefunction and also the base distribution of the proton flow model. Variational free energy calculation finetunes the networks and provides a solution for the dense hydrogen system.
We aim to study the equation of state of the hydrogen system near the melting line [27, 28], where protons are treated as classical point particles, and electrons are assumed to stay in the instantaneous ground state for any given proton configuration. To achieve this goal, we employ the variational free energy approach [35, 36, 29]. The corresponding objective function is given by
$F=\mathop{\mathbb{E}}_{\bm{X}\sim p(\bm{X})}[k_{B}T\ln p(\bm{X})+E(\bm{X})]\,,$ (2)
with $p(\mathbf{X})$ the proton Boltzmann distribution and $E(\bm{X})$ the internal energy of the system for given proton coordinates $\bm{X}$.
In the variational free energy method, the proton Boltzmann distribution is parameterized by a normalizing flow network [29]
$p(\bm{X})=\frac{e^{-E_{\mathrm{base}}(\bm{Z})/k_{B}T}}{\mathcal{Z}_{\mathrm{base}}}\left|\det{\left(\frac{\partial\bm{Z}}{\partial\bm{X}}\right)}\right|\,,$ (3)
where $E_{\mathrm{base}}$ is the potential energy surface (PES) provided by the base model, $\bm{Z}$ are a set of collective variables that are transformed from the proton coordinates $\bm{X}$ via a neural network with a non-degenerate Jacobian. By parametrizing the proton Boltzmann distribution in this way, the normalizing flow model learns to correct the potential bias in the base PES. The partition function of the base model, $\mathcal{Z}_{\mathrm{base}}$, acts as a constant factor across different parameters and does not affect the optimization process. The Jacobian determinant of the transformation in Eq. (3) ensures that the normalization factor of the proton distribution remains independent of the flow model parameters. With a perfectly optimized flow model, the probability distribution will converge to the target distribution $p(\bm{X})\propto e^{-E(\bm{X})/k_{B}T}$.
To avoid systematic bias and maintain a fully variational calculation framework, we compute the PES by parametrizing a trial wave function with neural networks, similar to the variational Monte Carlo (VMC) method. The ground-state energy is defined as
$E(\bm{X})=\mathop{\mathbb{E}}_{\bm{R}\sim|\Psi(\bm{R};\bm{X})|^{2}}\left[\frac{H\Psi(\bm{R};\bm{X})}{\Psi(\bm{R};\bm{X})}\right]\,,$ (4)
where $\Psi(\bm{R};\bm{X})$ denotes the ground-state electronic wave function for a given proton configuration $\bm{X}$. It is noteworthy that we use an ensemble VMC approach, as described by Xie et al. [29], in which the wave function is universally parameterized and jointly optimized for any proton coordinates sampled from the given distribution to achieve the ground state. Physical quantities such as energy, entropy, and pressure can be estimated from the samples generated by the proton Boltzmann distribution and electronic wavefunction after the free energy optimization.
To better leverage prior knowledge about the hydrogen system within the optimization framework, we design a two-stage optimization scheme as shown in Figure 1. First, proton configurations are generated through molecular dynamics sampling using DeePMD, a machine learning-based potential energy model [37]. Next, we pretrain the electronic wavefunction on the ensemble of proton configurations [38, 39]. Finally, the wavefunction and the flow model—where the DeePMD model serves as the base distribution [Eq. (3)]—are jointly optimized. This approach ensures that the variational free energy calculation serves as a fine-tuning stage that corrects possible biases in the DeePMD model for the proton distribution. In the following subsection, we will discuss general considerations for designing neural network architectures for variational free energy optimization of liquid hydrogen.
II.2 Neural networks
As shown in Eq. (3), the proton Boltzmann distribution is constructed using a flow model. A physically valid proton Boltzmann probability must be invariant under the translation of all protons, periodic transformations of individual protons, and permutations of protons. To ensure these symmetries, the flow model must itself be invariant under translations and periodic transformations, while also being equivariant under proton permutations. In practice, we achieve this by using a neural network that maps the proton coordinates $\bm{X}$ to a set of collective variables $\bm{Z}$ through an equivariant backflow transformation inspired by a Ferminet-like design [40]. For a comprehensive explanation of the flow network, readers are directed to Appendix A.2.
For for the electron wavefunction ansatz, we design a modified Ferminet [40] / Deepsolid [41] network architecture to satisfy the required symmetry constraints in the liquid phase. The translation invariance of the electron wavefunction under translations of both the electrons and protons requires $\Psi(\bm{R};\bm{X})=\Psi(\bm{R}+\bm{s};\bm{X}+\bm{s})$, where $\bm{s}$ is any constant vector in $\mathbb{R}^{3}$. Additionally, we impose the special twisted boundary conditions (TBC) [42] on the electronic degrees of freedom to mitigate finite-size effects. Under these conditions, wrapping an electron around the simulation cell introduces an additive phase change to the wavefunction, while the wavefunction remains periodic with respect to the proton degrees of freedom
$\Psi(\dots,\bm{r}_{i}+\bm{n},\dots;\bm{X})=e^{2\pi i\bm{w}\cdot\bm{n}}\Psi(\dots,\bm{r}_{i},\dots;\bm{X}),$ (5)
$\Psi(\bm{R};\cdots,\bm{x}_{I}+\bm{n},\cdots)=\Psi(\bm{R};\cdots,\bm{x}_{I},\cdots),$ (6)
with $\bm{n}=n_{1}\bm{a}_{1}+n_{2}\bm{a}_{2}+n_{3}\bm{a}_{3}$, $n_{\mu}\in\mathbb{Z}$ and $\mu=1,2,3$, and $\bm{w}$ the specified twist of the wavefunction. Furthermore, the wavefunction should be symmetric (anti-symmetric) with respect to the permutation of protons (electrons of the same spin):
$\Psi(\bm{R};\sigma(\bm{X}))=\Psi(\bm{R};\bm{X})\,,$ (7)
$\Psi(\sigma(\bm{R});\bm{X})=(-1)^{|\sigma|}\Psi(\bm{R};\bm{X})\,.$ (8)
The two major improvements we made compared to the network architectures in Ref. 29 are as follows: First, we introduce a multi-scale transformation for constructing periodic features, which enhances the feature representation in both the flow and electron wavefunction neural networks. Detailed explanations can be found in Appendix Section A.1. Second, we incorporate a gated attention layer within the message-passing process of the electron wave function ansatz. This design improves the network’s generalization ability, as elaborated in Appendix A.3.
III Results
In this section, we first present benchmark results to demonstrate the accuracy of the electron wavefunction, then we move on to physical results.
III.1 Benchmarks
The variational free energy calculations at different pressures utilize a neural network wavefunction with the same architecture. Therefore, we aim to demonstrate the flexibility of the electron wavefunction in this section, so that the calculation will not be biased towards one particular phase during the free energy optimization procedure.
In our framework, the wavefunction is pretrained using ensemble VMC training. In the pretraining procedure, the proton configurations are sampled using DeePMD simulations on a grid within the temperature-$r_{s}$ thermodynamic state space. The temperature ranges from 1000 to 2000 K in intervals of 200 K, while $r_{s}$ ranges from 1.35 to 1.50 in intervals of 0.02. A total of 6400 configurations are generated, and a random subset with 512 configurations is used for each training step. The batch size for electron configurations is set to 8.
Table 1: MAE in ground-state energy of ensemble VMC training for systems containing 32 hydrogen atoms.
mHa/atom
1231000K 1231200K 1231400K
$r_{s}=1.40$ 1230.81 1231.16 1230.83
$r_{s}=1.44$ 1231.61 1231.63 1231.63
To test the transferability of our wavefunction, we benchmark its performance by comparing the ground-state energies predicted by the multi-configuration trained model against those of the single-configuration trained model. Table 1 reports the mean absolute error (MAE) calculated using three test configurations for $r_{s}=1.40$ and $r_{s}=1.44$ at three different temperatures. These tested thermodynamic states are not included in the training ensemble. In all cases, the MAE remains below 2 mHa per atom, demonstrating that our pre-trained wavefunction is able to reliably provides accurate energy estimates across an ensemble of configurations, thus satisfying the requirements for variational free energy training.
Figure 2: The (a) pressure and (b) entropy per particle versus density parameteter computed at $T=1200$K for $N=32$ hydrogen atoms. The variational free energy calculations reveal significant features due to the formation of a crystalline structure at low densities.
Figure 3: The proton-proton radial distribution function of the DeePMD base and the deep variational free energy prediction computed at $T=1200$K.
Figure 4: (a) The radius distributuion function of protons computed at $T=1200$K and $r_{s}=1.43$ along with the one after quench. (b) A snapshot of the protons shown as white spheres. The red spheres denote the center of mass of hydrogen molecules. (c) The simulated XRD spectra for an ensemble of configurations obtained at 1200K.
III.2 High-temperature solids
In this subsection, we present the results of variational free energy optimization for the dense hydrogen system at 1200 K, focusing on the emergence of crystalline order.
Figure 2 shows the pressure and entropy as a function of the density parameter $r_{s}$ at $T=1200$K. The change in slope at $r_{s}=1.42$ in the pressure curve indicates a discontinuous phase transition. Since the value of the partition function of the base distribution enters the measurement of entropy (see Eq. (3)), we performed annealed importance sampling [43] on the DeePMD model to compute its free energy $-\ln\mathcal{Z}_{DP}$ in order to obtain the entropy of the system. The entropy curves shown in panel (b) provides further insight into the two states, with the sharp decrease in entropy at $r_{s}=1.42$ suggesting a transition to a more ordered state.
To further analyze the phase transition, we plot the proton-proton radial distribution function (RDF) of the trained model alongside the DeePMD base distribution in Figure 3. Compared to the uniform base distribution employed in Ref. 29, the DeePMD model provides a more informative reference for the proton distribution. However, around the critical point at $r_{s}=1.42$, the variational free energy optimization results in noticeable deviations in the RDF. The transition into a molecular state is characterized by the emergence of the molecular peak at around $1.4~{}a_{B}$ as the $r_{s}$ value increases. Additionally, the presence of extra features around $2~{}a_{B}$ and $3~{}a_{B}$ suggests that the system may develop additional structures other than a molecular liquid.
We perform a detailed analysis of the proton configurations for $r_{s}$ values larger than 1.41 to gain deeper insight into the molecular phase. Panel (a) of Figure 4 presents a snapshot of the sampled atomic configurations at $r_{s}=1.43$. Hydrogen molecules are identified by pairing the closest hydrogen atoms, with their centers of mass marked by red spheres. Notably, these red spheres do not follow a random liquid arrangement, but instead form an ordered structure reminiscent of a solid. The averaged simulated X-ray diffraction (XRD) spectra for an ensemble of hydrogen configurations at three different $r_{s}$ values are shown in Figure 4 (c). The XRD spectra of the structures do not show significant similarities to the $Fmmm$ structure reported in Ref. 27.
To further investigate the crystal symmetry of the discovered solid structure, we quench the atomic configurations by sampling from $p(\bm{X})^{\beta}$ with $\beta=20$. Figure  4 (b) shows the RDF of the original and quenched structures. As the system cools down, the features around $2~{}a_{B}$ and $3~{}a_{B}$ become more prominent. Symmetry analysis of the centers of mass of the hydrogen atoms in the quenched samples reveals that most of the configuraionts belong to the $R\bar{3}c$ space group. On the other hand, symmetry analysis of the hydrogen atom coordinates reveals predominantly $P1$ symmetry, indicating a lack of orientational order in the H2 molecules.
IV Conclusion and Discussion
In this work, we observe evidence of a direct transition from atomic liquid to molecular crystal within a temperature regime believed to be relevant to the liquid-liquid transition in the dense hydrogen system. This finding arises from a variational free energy calculation with many-body precision. Moreover, our results demonstrate that using many-body precision energy, the flow model can transform the structureless liquid state of the base proton distribution into a solid-state solution. However, since the simulation is performed with a finite box containing 32 hydrogen atoms, finite-size effects may bias both the existence and the structure of the discovered solid states. Nevertheless, this finding aligns with recent discoveries of a possible solid-state phase in similar temperature and pressure regimes obtained through different computational methods and much larger simulation cells [27, 28], although with different detailed characterizations of the solid-state phase. Collectively, these works suggest that the liquid to solid transition may ocurr at a higher temperature than previously believed. Replacing the expensive wavefunction model with a cheaper yet accurate machine learning force field could allow larger scale calculations, potentially bypassing some of the limitations of this study. Furthermore, nuclear quantum effects can be incorporated into the calculation by generalizing the proton distribution to a variational density matrix, using the neural canonical transformation approach [44, 45].
Acknowledgements.
We thank Guglielmo Mazzola, Yubo Yang, and Qi Zhang for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. 12122103, No. 92270107, No. T2225018, No. 12188101, No. T2121001 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.
Appendix A Details of the neural network models
A.1 periodic features
The first step in constructing the neural networks for flow and the electron wavefunction is to introduce a periodic function $\bm{f}$ that transforms the coordinates of protons and electrons into a form that remains invariant under both spatial translations of all particles by any vector $\bm{s}\in\mathbb{R}^{3}$ and the periodic transformation of any individual particle by the lattice vector $\bm{n}$, i.e.
$\bm{f}(\bm{\Xi})=\bm{f}(\bm{\Xi}+\bm{s}),$ (9)
$\bm{f}(\dots,\bm{\xi}_{\alpha},\dots)=\bm{f}(\dots,\bm{\xi}_{\alpha}+\bm{n},\dots).$ (10)
We use the following periodic feature in our networks
$\bm{f}_{\alpha\beta}(\bm{\Xi})=\mathrm{concat}\Big{(}\bm{f}^{\mathrm{ds}}_{\alpha\beta},\>\cos(\pi\bm{\sigma}_{\alpha\beta})\Big{)}\,,$ (11)
in which $\bm{\sigma}_{\alpha\beta}=\sum_{\mu}\sigma_{\alpha\beta}^{\mu}\bm{a}_{\mu}$, $\sigma_{\alpha\beta}^{\mu}=\hat{\sigma}_{\alpha\beta}^{\mu}-2\lfloor(\hat{\sigma}_{\alpha\beta}^{\mu}+1)/2\rfloor$ and $\hat{\sigma}_{\alpha\beta}^{\mu}=2\bm{\xi}_{\alpha\beta}\cdot\bm{a}^{\ast}_{\mu}$. Here $\hat{\sigma}_{\alpha\beta}^{\mu}$ represents the direct coordinates of the relative position $\bm{\xi}_{\alpha\beta}=\bm{\xi}_{\alpha}-\bm{\xi}_{\beta}$, while $\sigma_{\alpha\beta}^{\mu}$ normalizes these coordinates to the range $[-1,1]$ by applying a shift. The $\cos$ operator is applied component-wise to the vector $\bm{\sigma}_{\alpha\beta}$ in the latter part, and the term $\bm{f}^{\mathrm{ds}}_{\alpha\beta}$ in this expression is the periodic features introduced by the deep-solid network [41]
$\bm{f}^{\mathrm{ds}}_{\alpha\beta}(\bm{\Xi})=\mathrm{concat}\Big{(}\frac{1}{2\pi}\sum_{\mu=1}^{3}g(\sigma_{\alpha\beta}^{\mu})\bm{a}_{\mu},\>d(\bm{\sigma}_{\alpha\beta})\Big{)}\,,$ (12)
in which
$d(\bm{\sigma})^{2}=\frac{1}{4\pi^{2}}\sum_{\mu\nu}M_{\mu\nu}(\bm{\sigma})\>\bm{a}_{\mu}\cdot\bm{a}_{\nu}\,,$ (13)
$M_{\mu\nu}(\bm{\sigma})=f^{2}(\sigma^{\mu})\delta_{\mu\nu}+g(\sigma^{\mu})g(\sigma^{\nu})(1-\delta_{\mu\nu})\,,$ (14)
$f(\sigma)=|\sigma|\Big{(}1-\frac{|\sigma|^{3}}{4}\Big{)}\,,\quad g(\sigma)=\sigma\Big{(}1-\frac{3}{2}|\sigma|+\frac{1}{2}|\sigma|^{2}\Big{)}\,.$ (15)
It is straightforward to verify that the features constructed via Eq. (11) satisfy the conditions specified in Eqs. (9) and (10).
To enhance the feature representation, we introduce a multi-scale transformation for the deep-solid features,
$\mathrm{multi\textunderscore scale}(\,\bm{f}^{\mathrm{ds}}_{\alpha\beta}\,)=\mathrm{concat}\Big{(}\{\,s_{i}\cdot\bm{f}^{\mathrm{ds}}_{\alpha\beta}\,\}_{i\in S}\Big{)}\,,$ (16)
where $S$ is a set of all possible scales. For instance, when considering 16 protons, $S=\{1,2,4\}$. The multi-scale transformation multiplies the deep-solid feature by each scale in the set $S$ and concatenates the results. The multi-scale transformed feature reads
$\bm{f}^{\mathrm{ms}}_{\alpha\beta}(\bm{\Xi})=\mathrm{concat}\Big{(}\mathrm{multi\textunderscore scale}(\bm{f}^{\mathrm{ds}}_{\alpha\beta}),\>\cos(\pi\bm{\sigma}_{\alpha\beta})\Big{)}.$ (17)
This feature is employed by default throughout this work unless stated otherwise.
A.2 Flow network
Algorithm 1 Flow ansätz
1:Proton coordinates $\bm{x}_{I}$.
2:$\bm{x}_{IJ}\leftarrow\bm{x}_{I}-\bm{x}_{J}$
3:$\bm{f}^{\mathrm{ms}}_{IJ}\leftarrow\textrm{features}(\bm{x}_{IJ})$triangleright Eq. (17)
4:$\bm{h}^{(0)}_{I}\leftarrow{\mathrm{linear}}\left(\frac{1}{N}\sum_{J}\bm{f}^{\mathrm{ms}}_{IJ}\right)\quad\bm{h}^{(0)}_{IJ}\leftarrow{\mathrm{linear}}\left(\bm{f}^{\mathrm{ms}}_{IJ}\right)$
5:for $i=0$ to $L-1$ do
6:     $\bm{g}^{(l)}_{I}\leftarrow\textrm{concat}\left(\bm{h}^{(l)}_{I},\frac{1}{N}\sum_{I}\bm{h}^{(l)}_{I},\frac{1}{N}\sum_{J}\bm{h}^{(l)}_{IJ}\hat{\bm{h}}^{(l)}_{J}\right)$
7:     $\bm{h}^{(l+1)}_{I}\leftarrow\frac{1}{\sqrt{2}}\bm{h}^{(l)}_{I}+\frac{1}{\sqrt{2}}\textrm{dense\textunderscore layer}\big{(}\bm{g}^{(l)}_{I}\big{)}$
8:     $\bm{h}^{(l+1)}_{IJ}\leftarrow\frac{1}{\sqrt{2}}\bm{h}^{(l)}_{IJ}+\frac{1}{\sqrt{2}}\textrm{dense\textunderscore layer}\big{(}\bm{h}^{(l)}_{IJ}\big{)}$
9:end for
10:$\bm{h}_{I}(\bm{X})\leftarrow\bm{h}^{(L)}_{I}$
11:$\bm{Z}_{I}(\bm{X})\leftarrow\bm{X}_{I}+{\text{linear}}\,(\bm{h}_{I}\big{(}\bm{X})\big{)}$
12:return $\bm{Z}_{I}(\bm{X})$
In Algorithm 1, we introduce single-particle features $\bm{h}_{I}$ and pair-particle features $\bm{h}_{IJ}$ with dimensions ${d^{p}_{1}}$ and ${d^{p}_{2}}$. These features are initialized using a multi-scale periodic feature function (line 2) and trainable linear mappings (line 3) which reshape their feature dimensions to ${d^{p}_{1}}$ and ${d^{p}_{2}}$. The features are then subsequently updated through $L$ message-passing layers, with the update at the $(l+1)$-th layer described by lines 5-7. The hat over $\bm{h}^{(l)}_{J}$ indicates that the single-particle feature is linearly projected to the dimension $d^{p}_{2}$, so the multiplication is performed component-wise. The “dense_layer” operation consists of a trainable linear mapping followed by an element-wise hyperbolic tangent activation. Finally, a trainable linear mapping is applied to the feature dimension $d^{p}_{1}$ to produce a 3-dimensional vector, and the coordinates are updated via a backflow architecture (line 10).
A.3 Electron wavefunction
Algorithm 2 Wavefunction ansätz
1:Coordinates $\bm{\xi}_{\alpha}$, one-hot $\bm{t}_{\alpha}$, kpoints $\bm{k}_{m}$, twist $\bm{w}$, scale $r_{s}$.
2:$\bm{\phi}_{i}^{\delta},\bm{h}_{\alpha}\leftarrow\textrm{orbitals}(\bm{\xi}_{\alpha},\bm{t}_{\alpha},\bm{w},r_{s})$triangleright Algorithm 3
3:$\bm{e}_{i}^{\delta}\leftarrow\textrm{envelope}(\bm{h}_{I},\bm{r}_{iI},r_{s})$triangleright Eq. (20) and (21)
4:$\bm{\phi}_{i}^{\delta}\leftarrow\bm{e}_{i}^{\delta}\odot\bm{\phi}_{i}^{\delta}$
5:$D^{\delta}_{ij}\leftarrow\sum_{\mu}W_{\mu}\phi^{\delta}_{i,\mu}\phi^{\delta}_{j,\mu}$triangleright $i\in\uparrow$, $j\in\downarrow$
6:$\lambda_{m}^{\delta}\leftarrow\textrm{MLP}(\bm{w},r_{s})$
7:$\hat{\bm{r}}_{i}\leftarrow\bm{r}_{i}+\textrm{linear}(\bm{h}_{i})$
8:$E^{\delta}_{ij}\leftarrow\frac{1}{\sqrt{n_{k}V}}\sum_{m=1}^{n_{k}}\lambda_{m}^{\delta}e^{2\pi i(\bm{k}_{m}+\bm{w})\cdot(\hat{\bm{r}}_{i}-\hat{\bm{r}}_{j})}$triangleright $i\in\uparrow$, $j\in\downarrow$
9:$\Psi\leftarrow\sum_{\delta}\det(E^{\delta}\odot D^{\delta})$triangleright $\odot$ element-wise multiplication
10:return $\Psi$
The wavefunction ansatz is outlined in Algorithm 2. In this algorithm, electrons and protons are treated equivalently and are distinguished by a one-hot type coding $\bm{t}_{\alpha}$, i.e. electrons in the $\uparrow$ and $\downarrow$ channels are encoded as $(1,0,0)$ and $(0,1,0)$, respectively, while protons are encoded as $(0,0,1)$. This type coding can be extended to systems with heavier nucleis. In addition to the coordinates and type coding, we also input the k-space sampling points $\bm{k}_{m}$, the twist $\bm{w}$, and the scale $r_{s}$ as parameters to the wavefunction. With these inputs, the coordinates are first processed to construct the feature function $\bm{h}_{\alpha}$, the electron orbital $\bm{\phi}_{i}^{\delta}$, and the electron envelope $\bm{e}_{i}^{\delta}$ (lines 1-2). The geminal matrix [46] $D_{ij}^{\delta}$ is computed as a bilinear form of the spin-up and spin-down electron orbitals, with a diagonal trainable parameter matrix $W_{\mu}$ (line 4). The geminal envelope $E_{ij}^{\delta}$ is constructed as a linear combination of plane waves with twist boundary conditions and a backflow transformation of the electron coordinates (lines 5-7). Finally, the wavefunction is expressed as a linear combination of the determinants of the geminals (line 10).
Algorithm 3 Electron orbitals
1: Coordinates $\bm{\xi}_{\alpha}$, one-hot $\bm{t}_{\alpha}$, twist $\bm{w}$ and scale $r_{s}$.
2:$\bm{\xi}_{\alpha\beta}\leftarrow\bm{\xi}_{\alpha}-\bm{\xi}_{\beta}$
3:$\bm{f}^{\mathrm{ms}}_{\alpha\beta}\leftarrow\textrm{features}(\bm{\xi}_{\alpha\beta})$triangleright Eq. (17)
4:$\bm{v}\leftarrow\mathrm{concat}(\bm{w},r_{s})$,
5:$\bm{h}^{(0)}_{\alpha}\leftarrow\mathrm{linear}\Big{(}\mathrm{concat}\Big{(}\frac{1}{n}\sum_{\beta}\bm{f}^{\mathrm{ms}}_{\alpha\beta},\,\bm{t}_{\alpha},\bm{v}\Big{)}\Big{)}$,
6:$\bm{h}^{(0)}_{\alpha\beta}\leftarrow\mathrm{linear}\Big{(}\mathrm{concat}\Big{(}\bm{f}^{\mathrm{ms}}_{\alpha\beta},\bm{t}_{\alpha},\bm{t}_{\beta},\bm{v}\Big{)}\Big{)}$,
7:for $l=0$ to $L-1$ do
8:     $\bm{g}^{(l)}_{\alpha}\leftarrow\mathrm{intermediate\textunderscore feature}(\bm{h}^{(l)}_{\alpha})$triangleright see Eq. (18) and (19)
9:     $\bm{h}^{(l+1)}_{\alpha}\leftarrow\frac{1}{\sqrt{2}}\bm{h}^{(l)}_{\alpha}+\frac{1}{\sqrt{2}}\textrm{dense\textunderscore layer}\big{(}\bm{g}^{(l)}_{\alpha}\big{)}$
10:     $\bm{h}^{(l+1)}_{\alpha\beta}\leftarrow\frac{1}{\sqrt{2}}\bm{h}^{(l)}_{\alpha\beta}+\frac{1}{\sqrt{2}}\textrm{dense\textunderscore layer}\big{(}\bm{h}^{(l)}_{\alpha\beta}\big{)}$
11:     if do_attn then
12:         $\bm{h}^{(l+1)}_{\alpha\beta}\leftarrow\frac{1}{\sqrt{2}}\bm{h}^{(l)}_{\alpha\beta}+\frac{1}{\sqrt{2}}\textrm{gated\textunderscore atten\textunderscore layer}\big{(}\bm{h}^{(l)}_{\alpha\beta}\big{)}$
13:     end if
14:end for
15:$\bm{\phi}^{\delta}_{i}\leftarrow\textrm{linear}(\bm{h}_{i}^{(L)})$triangleright $\delta$ is the index of determinants
16:return $\bm{\phi}^{\delta}_{i}$, $\bm{h}_{\alpha}^{(L)}$
The electron orbitals are constructed in a Ferminet-like manner, with single- and pair-particle features defined symmetrically for both electrons and protons. In each layer, an intermediate feature $\bm{g}$ is constructed as follows
$\bm{g}^{(l)}_{I}=\textrm{concat}\bigg{(}$ $\bm{h}^{(l)}_{I},\frac{1}{N}\sum_{I}\bm{h}^{(l)}_{I},\frac{1}{N}\sum_{I}\bm{h}^{(l)}_{I},\frac{1}{n^{\uparrow}}\sum_{i\in\uparrow}\bm{h}^{(l)}_{Ii}\hat{\bm{h}}^{(l)}_{i},$
$\frac{1}{n^{\downarrow}}\sum_{i\in\downarrow}\bm{h}^{(l)}_{Ii}\hat{\bm{h}}^{(l)}_{i},\frac{1}{N}\sum_{J}\bm{h}^{(l)}_{IJ}\hat{\bm{h}}^{(l)}_{J}\bigg{)},$ (18)
$\bm{g}^{(l)}_{i}=\textrm{concat}\bigg{(}$ $\bm{h}^{(l)}_{i},\frac{1}{n^{\uparrow}}\sum_{i}\bm{h}^{(l)}_{i},\frac{1}{n^{\downarrow}}\sum_{i}\bm{h}^{(l)}_{i},\frac{1}{n^{\uparrow}}\sum_{j\in\uparrow}\bm{h}^{(l)}_{ij}\hat{\bm{h}}^{(l)}_{j},$
$\frac{1}{n^{\downarrow}}\sum_{j\in\downarrow}\bm{h}^{(l)}_{ij}\hat{\bm{h}}^{(l)}_{j},\frac{1}{N}\sum_{I}\bm{h}^{(l)}_{iI}\hat{\bm{h}}^{(l)}_{I}\bigg{)}.$ (19)
The hat over $\bm{h}^{(l)}_{\alpha}$ indicates that the single-particle feature is linearly projected to the same dimension as $\bm{h}_{\alpha\beta}$, enabling component-wise multiplication.
We introduce a gated attention layer during the message-passing procedure to enhance the network’s generalization ability. The gated self-attention Algorithm 4 is like the standard self-attention algorithm, except the gate term inspired by the AlphaFold [47]. The query, key and value are the linear mappings from the pair-particle channel. The attention map is formed by the softmax applied to the inner product of queries and keys connecting to the same particle. The gate is applied as a component-wise multiplication on the attention mapped values. Finally, the values in different heads are concatenated and linearly mapped to give the output pair-particle channel.
Algorithm 4 Gated self-attention
1:Pair-particle channel $\bm{h}_{\alpha\beta}$
2:$\bm{h}_{\alpha\beta}\leftarrow\textrm{layer\textunderscore norm}(\bm{h}_{\alpha\beta})$
3:$\bm{q}^{h}_{\alpha\beta},\bm{k}^{h}_{\alpha\beta},\bm{v}^{h}_{\alpha\beta}\leftarrow\textrm{linear\textunderscore no\textunderscore bias}(\bm{h}_{\alpha\beta})$
4:$\bm{g}^{h}_{\alpha\beta}\leftarrow\mathrm{sigmoid}\left(\textrm{linear}(\bm{h}_{\alpha\beta})\right)$
5:$m^{h}_{\alpha\gamma\beta}\leftarrow\textrm{softmax}_{\gamma}(\bm{q}^{h}_{\alpha\beta}{}^{\top}\bm{k}^{h}_{\gamma\beta}/\sqrt{d_{h}})$
6:$\bm{u}^{h}_{\alpha\beta}\leftarrow\bm{g}^{h}_{\alpha\beta}\odot\sum_{\gamma}m^{h}_{\alpha\gamma\beta}\bm{v}^{h}_{\gamma\beta}$triangleright $\odot$ element-wise multiplication
7:$\hat{\bm{h}}_{\alpha\beta}\leftarrow\textrm{linear\textunderscore no\textunderscore bias}(\textrm{concat}_{h}(\bm{u}^{h}_{\alpha\beta}))$
8:return $\hat{\bm{h}}_{\alpha\beta}$
Figure 5: Comparison of ensemble training curves with and without the attention layer.
Figure 5 shows a comparison of the ensemble VMC training results for systems with 32 hydrogen atoms, both with and without the attention layer. The training curve clearly shows that the attention layer significantly enhances the expressibility of the wavefunction in the multi-configuration case, leading to a lower variational energy.
The orbital envelope is given by:
$\bm{e}^{\delta}_{i}=\sum_{I}\bm{\pi}^{\delta}_{I}\odot\exp(-d_{iI}\bm{\Sigma}^{\delta}_{I})$ (20)
$\bm{\pi}^{\delta}_{I}=1.0+\textrm{MLP}(\bm{h}_{I}),\quad\bm{\Sigma}^{\delta}_{I}=r_{s}+\textrm{MLP}(\bm{h}_{I}),$ (21)
where the periodic distance $d_{iI}=d(\bm{\sigma}_{iI})$ (defined by Eq. (13)) between electron $i$ and proton $I$ is scaled by the trainable parameters $\bm{\Sigma}^{\delta}_{I}$ to provide the input for the exponent. The exponent is then scaled by the trainable parameters $\bm{\pi}^{\delta}_{I}$ to form the envelope of proton $I$. The envelopes are summed over all protons to provide the contribution of the $i$-th orbital of determinant $\delta$. Notably, all electrons share the same parameters $\bm{\pi}^{\delta}_{I}$ and $\bm{\Sigma}^{\delta}_{I}$. The envelope is isotropic, meaning that the decay with respect to the distance $d$ is uniform across all three spatial dimensions.